OpenStudy (anonymous):

Factor each trinomial below. Please show your work and check your answer. 1)a2 – 5a – 20

OpenStudy (anonymous):

Is your question right?? Because it is not factorisable...

OpenStudy (anonymous):

well thats what it said but i hered i might have to use quuadratic formula

OpenStudy (anonymous):

Yeah I am thinking the same.. \[a = \frac{5 \pm \sqrt{25 + 80}}{2} \implies a = \frac{5 \pm \sqrt{105}}{2} = ??\] This is not factorisable I think..